### v1

# Rephasing invariant formulae for general parameterizations of flavor mixing matrix and exact sum rules with unitarity triangles

Abstract
In this letter, we present rephasing invariant formulae $\delta^{(\alpha i)}=\arg[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}]$ for CP phases $\delta^{(\alpha i)}$ associated with nine Euler-angle-like parameterizations of a flavor mixing matrix. Here, $\alpha$ and $i$ denote the row and column carrying the trivial phases in a given parameterization. Furthermore, we show that the phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of unitarity triangles satisfy compact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}$ and $\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ where all indices are taken cyclically modulo three. These twelve relations are natural generalizations of the previous result $\delta_{\mathrm{PDG}}+\delta_{\mathrm{KM}}=\pi-\alpha+\gamma.$
preprint: STUPP-25-285
I Introduction
The flavor mixing of quarks and leptons has been one of the essential pieces for the development of particle physics, and it is also an intrinsically interesting subject of mathematical investigation. Among various representations of the mixing matrix Kobayashi:1973fv ; Chau:1984fp ; Hall:1993ni ; Kaur:2023ypg ; Yang:2020qsa ; Yang:2020goc ; Yang:2024ulq ; Yang:2025yst , nine Euler-angle-like parameterizations in terms of three rotation matrices and a distinct CP phase admit nine possible forms Fritzsch:1995nx ; Rasin:1997pn . However, at first glance, these nine phases seem to take chaotic values, showing no apparent pattern.
On the other hand, unitarity conditions of the mixing matrix give rise to six unitarity triangles Wu:1994di ; Xing:2009eg ; Harrison:2009bz ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Harrison:2025rkp and relations between the CP phases and the unitarity triangles has also been extensively studied Koide:2004gj ; Koide:2008yu ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2011bq ; Zhou:2011xm ; Qin:2011ub ; Qin:2010hn ; Zhang:2012ys ; Li:2012zxa ; Zhang:2012bk . These nine angles also seem to have independent values, showing no obvious structure.
More recently, rephasing invariant formulae for the CP phases Yang:2025hex ; Yang:2025law ; Yang:2025ftl produce a simple sum rule relating the CP phases and angles $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$. This led to a discovery of an underlying order between the random CP phases and angles.
In this letter, to establish exact sum rules connecting the nine CP phases and the nine angles of unitarity triangles, we first derive rephasing invariant formulae for each CP phase in terms of elements of the mixing matrix $V$ and $\det V$. Next, we present sum rules between the phases and angles, showing that the previously known sum rule arises as a special case of these relations.
II Rephasing invariant formulae for CP phases in general parameterizations of flavor mixing matrix
In this section, following our previous works Yang:2025hex ; Yang:2025law ; Yang:2025ftl , we derive rephasing invariant formulae for CP phases in Euler-angle-like parameterizations of the flavor mixing matrix $V$. For this purpose, we adopt the nine parameterizations proposed by Fritzsch and Xing Fritzsch:1997st .
We begin by defining the $2\times 2$ rotation matrix $R_{ij}$ as follows,
$R_{12}(\theta)=\begin{pmatrix}c_{\theta}&s_{\theta}&0\\ -s_{\theta}&c_{\theta}&0\\ 0&0&1\\ \end{pmatrix},~~~R_{23}(\sigma)=\begin{pmatrix}1&0&0\cr 0&c_{\sigma}&s_{\sigma}\\ 0&-s_{\sigma}&c_{\sigma}\\ \end{pmatrix},~~R_{31}(\tau)=\begin{pmatrix}c_{\tau}&0&s_{\tau}\\ 0&1&0\\ -s_{\tau}&0&c_{\tau}\\ \end{pmatrix}$, (1)
where $s_{\theta}=\sin\theta,c_{\theta}=\cos\theta.$ Furthermore, we define complex rotation matrices $R_{12}(\theta,\delta)$, $R_{23}(\sigma,\delta)$, and $R_{31}(\tau,\delta)$ by replacing $1\rightarrow e^{-i\delta}$ in Eq. (1). With these matrices, the mixing matrix $V$ admits nine different parameterizations Fritzsch:1997st
$P1:~V^{(33)}$ $=R_{12}(\theta)R_{23}(\sigma,\delta^{(33)})R^{-1}_{12}(\theta^{\prime})\,,~~P2:~V^{(11)}=R_{23}(\sigma)R_{12}(\theta,\delta^{(11)})R^{-1}_{23}(\sigma^{\prime})\,$, (2)
$P3:~V^{(13)}$ $=R_{23}(\sigma)R_{31}(\tau,\delta^{(13)})R_{12}(\theta)\,,~~~P4:~V^{(31)}=R_{12}(\theta)R_{31}(\tau,\delta^{(31)})R^{-1}_{23}(\sigma)\,$, (3)
$P5:~V^{(22)}$ $=R_{31}(\tau)R_{12}(\theta,\delta^{(22)})R^{-1}_{31}(\tau^{\prime})\,,~~P6:~V^{(32)}=R_{12}(\theta)R_{23}(\sigma,\delta^{(32)})R_{31}(\tau)\,$, (4)
$P7:~V^{(12)}$ $=R_{23}(\sigma)R_{12}(\theta,\delta^{(12)})R^{-1}_{31}(\tau)\,,~~P8:~V^{(21)}=R_{31}(\tau)R_{12}(\theta,\delta^{(21)})R_{23}(\sigma)\,$, (5)
$P9:~V^{(23)}$ $=R_{31}(\tau)R_{23}(\sigma,\delta^{(23)})R^{-1}_{12}(\theta)\,.$ (6)
Here, in each of these parameterizations, a row $\alpha$ and a column $i$ have trivial phases, $V_{\alpha j},V_{\beta i}\in\mathbb{R}$. The notation $V^{(\alpha i)}$ and $\delta^{(\alpha i)}$ distinguish the corresponding matrix and CP phase. More explicitly, the conditions are
$\arg V_{\alpha 1}^{(\alpha i)}=\arg V_{\alpha 2}^{(\alpha i)}=\arg V_{\alpha 3}^{(\alpha i)}=\arg V_{1i}^{(\alpha i)}=\arg V_{2i}^{(\alpha i)}=\arg V_{3i}^{(\alpha i)}=0~{\rm or}~\pi\,.$ (7)
Since there is a duplication for the element $V_{\alpha i}^{(\alpha i)}$, there are effectively five independent conditions, and the phase $\pi$ appears either zero or two times. Finally, if the Majorana phases are omitted,
$\arg\det V^{(\alpha i)}=-\delta^{(\alpha i)}\,.$ (8)
In other words, the phase structure of these parameterizations $V^{(\alpha i)}$ is specified by these six conditions. Since the resulting CP phases are manifestly rephasing invariant, this construction remains valid even in the presence of Majorana phases for leptons.
Let us consider transforming a mixing matrix $V$ in an arbitrary basis into one of these parameterizations $V^{(\alpha i)}$. Suppose that unphysical phases are removed by general phase transformations:
$V^{(\alpha i)}=\Psi_{L}^{\dagger}V\Psi_{R},$ (9)
where $(\Psi_{L})_{\alpha\beta}=e^{i\gamma_{L\alpha}}\delta_{\alpha\beta}$ and $(\Psi_{R})_{ij}=e^{i\gamma_{Ri}}\delta_{ij}$ are diagonal phase matrices. The inverse transformation from $V^{(\alpha i)}$ to the original matrix $V$ is given by
$V=\Psi_{L}V^{(\alpha i)}\Psi_{R}^{\dagger}\,,~~V_{\beta j}=e^{i\gamma_{L\beta}}V^{(\alpha i)}_{\beta j}e^{-i\gamma_{Rj}}\,.$ (10)
From a relation $\arg\det V=-\delta^{(\alpha i)}+\sum_{\beta,j}(\gamma_{L\beta}-\gamma_{Rj})$ obtained by taking the argument of determinant, the CP phase $\delta^{(\alpha i)}$ is found to be
$\delta^{(\alpha i)}$ $=(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})-\arg\det V$ (11)
$=\arg[V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}]-\arg\det V=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right].$ (12)
In practice, since $V_{\alpha i}$ appears twice in the numerator, cancellations occur between the numerator and denominator:
$\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right]=\arg\left[\prod_{\beta\neq\alpha,j\neq i}{V_{\alpha j}V_{\beta i}\over V_{\alpha i}\det V}\right].$ (13)
Since six phases are required to eliminate unphysical phases $\gamma_{L\alpha}$ and $\gamma_{Ri}$, any other combinations have nontrivial phases of $V^{(\alpha i)}_{\beta j}$ explicitly.
Explicitly, the formulae are expressed as
$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$ (14)
$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$ (15)
$\delta^{(31)}$ $=\arg\left[{V_{32}V_{33}V_{11}V_{21}\over V_{31}\det V}\right],~~\delta^{(32)}=\arg\left[{V_{31}V_{33}V_{12}V_{22}\over V_{32}\det V}\right],~~\delta^{(33)}=\arg\left[{V_{31}V_{32}V_{13}V_{23}\over V_{33}\det V}\right].$ (16)
They provide a necessary and sufficient set of “fifth-order” invariants, because only nine such invariants can be constructed using $\det V$.
By using the CKM matrix parameters from the latest UTfit results UTfit:2022hsi
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$, (17)
$\sin\theta_{13}$ $=0.003714\pm 0.000092\,,~~~\delta=1.137\pm 0.022=65.15^{\circ}\pm 1.3^{\circ}\,$, (18)
the numerical evaluation at the best-fit values yields
$\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix}=\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.16^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}$. (19)
In particular, the standard PDG phase $\delta_{\rm PDG}$ and the original Kobayashi–Maskawa phase $\delta_{\rm KM}$ are given by
$\delta_{\rm PDG}=\delta^{(13)}\,,~~~\delta_{\rm KM}=\pi-\delta^{(11)}\,.$ (20)
Interestingly, a sum over any row or column of the phases $\delta^{(ij)}$ yields the same value
$\sum_{i}\delta^{(ij)}=\sum_{j}\delta^{(ij)}=\arg\left[{\prod_{\alpha,i}V_{\alpha i}\over\det V^{3}}\right]=273.49^{\circ}\,.$ (21)
For a cross check, we numerically compute $\cos\delta^{(ij)}$,
$\begin{pmatrix}\cos\delta^{(11)}&\cos\delta^{(12)}&\cos\delta^{(13)}\\ \cos\delta^{(21)}&\cos\delta^{(22)}&\cos\delta^{(23)}\\ \cos\delta^{(31)}&\cos\delta^{(32)}&\cos\delta^{(33)}\end{pmatrix}=\begin{pmatrix}-0.0425&-0.437&0.420\\ -0.924&0.9998&-0.421\\ 0.917&-0.917&-0.0603\\ \end{pmatrix}$. (22)
These values agree with values obtained from moduli of matrix elements,
$\cos\delta^{(11)}$ $=\frac{\left(1-|V_{11}|^{2}\right)^{2}|V_{22}|^{2}-|V_{11}V_{12}V_{21}|^{2}-|V_{13}V_{31}|^{2}}{2|V_{11}V_{12}V_{13}V_{21}V_{31}|}=-0.0425\,$, (23)
$\cos\delta^{(12)}$ $=\frac{\left(1-|V_{12}|^{2}\right)^{2}|V_{23}|^{2}-|V_{12}V_{13}V_{22}|^{2}-|V_{11}V_{32}|^{2}}{2|V_{11}V_{12}V_{13}V_{22}V_{32}|}=-0.437\,$, (24)
$\cos\delta^{(13)}$ $=-\frac{\left(1-|V_{13}|^{2}\right)^{2}|V_{22}|^{2}-|V_{12}V_{23}V_{13}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{12}V_{13}V_{23}V_{33}|}=0.420\,$, (25)
$\cos\delta^{(21)}$ $=-\frac{\left(1-|V_{21}|^{2}\right)^{2}|V_{12}|^{2}-|V_{11}V_{21}V_{22}|^{2}-|V_{23}V_{31}|^{2}}{2|V_{21}V_{22}V_{23}V_{11}V_{31}|}=-0.924\,$, (26)
$\cos\delta^{(22)}$ $=\frac{\left(1-|V_{22}|^{2}\right)^{2}|V_{33}|^{2}-|V_{22}V_{23}V_{32}|^{2}-|V_{12}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{12}V_{32}|}=0.9998\,$, (27)
$\cos\delta^{(23)}$ $=-\frac{\left(1-|V_{23}|^{2}\right)^{2}|V_{32}|^{2}-|V_{22}V_{23}V_{33}|^{2}-|V_{13}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{13}V_{33}|}=-0.421\,$, (28)
$\cos\delta^{(31)}$ $=-\frac{\left(1-|V_{31}|^{2}\right)^{2}|V_{22}|^{2}-|V_{21}V_{32}V_{31}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{21}V_{31}V_{32}V_{33}|}=0.917\,$, (29)
$\cos\delta^{(32)}$ $=-\frac{\left(1-|V_{32}|^{2}\right)^{2}|V_{23}|^{2}-|V_{22}V_{32}V_{33}|^{2}-|V_{12}V_{31}|^{2}}{2|V_{31}V_{32}V_{33}V_{12}V_{22}|}=-0.917\,$, (30)
$\cos\delta^{(33)}$ $=\frac{\left(1-|V_{33}|^{2}\right)^{2}|V_{11}|^{2}-|V_{13}V_{31}V_{33}|^{2}-|V_{23}V_{32}|^{2}}{2|V_{13}V_{31}V_{33}V_{23}V_{32}|}=-0.0603\,.$ (31)
Since $\delta^{(\alpha i)}$ also carries the sign of $\sin\delta^{(\alpha i)}$, it contains more information than an evaluation based on absolute values of matrix elements.
III Sum rules between CP phases in general parameterizations and unitarity triangles
It is noteworthy to revisit relations between the results and angles of unitarity triangles. The three angles of the unitarity triangle are rephasing invariants,
$\alpha=\arg\left[-{V_{td}V_{tb}^{*}\over V_{ud}V_{ub}^{*}}\right]=92.40^{\circ},~~\beta=\arg\left[-{V_{cd}V_{cb}^{*}\over V_{td}V_{tb}^{*}}\right]=22.49^{\circ},~~\gamma=\arg\left[-{V_{ud}V_{ub}^{*}\over V_{cd}V_{cb}^{*}}\right]=65.11^{\circ}.$ (32)
A general phase matrix $\Phi$ of these angles is defined in Harrison:2009bz . In particular, the elements in the second column of $\Phi$ correspond to $\beta$, $\alpha$, and $\gamma$, respectively.
$\Phi\equiv\begin{pmatrix}\arg(-\Pi_{ud}^{*})&\arg(-\Pi_{us}^{*})&\arg(-\Pi_{ub}^{*})\\ \arg(-\Pi_{cd}^{*})&\arg(-\Pi_{cs}^{*})&\arg(-\Pi_{cb}^{*})\\ \arg(-\Pi_{td}^{*})&\arg(-\Pi_{ts}^{*})&\arg(-\Pi_{tb}^{*})\\ \end{pmatrix}=\begin{pmatrix}1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ \end{pmatrix}$. (33)
Here, $\Pi_{\alpha i}\equiv V_{\beta j}V_{\beta k}^{*}V_{\gamma k}V_{\gamma j}^{*}$ and the indices are defined cyclically. For better readability, all flavor indices are rewritten by using numbers as
$\Pi=\begin{pmatrix}V_{33}V_{32}^{*}V_{22}V_{23}^{*}&V_{31}V_{33}^{*}V_{23}V_{21}^{*}&V_{32}V_{31}^{*}V_{21}V_{22}^{*}\\ V_{13}V_{12}^{*}V_{32}V_{33}^{*}&V_{11}V_{13}^{*}V_{33}V_{31}^{*}&V_{12}V_{11}^{*}V_{31}V_{32}^{*}\\ V_{23}V_{22}^{*}V_{12}V_{13}^{*}&V_{21}V_{23}^{*}V_{13}V_{11}^{*}&V_{22}V_{21}^{*}V_{11}V_{12}^{*}\end{pmatrix}$. (34)
The sum over each row and column of the matrix $\Phi$ equals $180^{\circ}$,
$\sum_{i}\Phi_{ij}=\sum_{j}\Phi_{ij}=\pi\,.$ (35)
Since solving general sum rules is a difficult task, we employ a transfer-matrix approach. Defining a phase matrix $\Delta$ and a transfer matrix $T$,
$\Delta\equiv\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix},~~T\equiv\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\\ \end{pmatrix}$, (36)
one finds the following relations
$\Delta T-\Delta=\begin{pmatrix}\arg\frac{V_{11}^{2}V_{22}V_{32}}{V_{12}^{2}V_{21}V_{31}}&\arg\frac{V_{12}^{2}V_{23}V_{33}}{V_{13}^{2}V_{22}V_{32}}&\arg\frac{V_{13}^{2}V_{21}V_{31}}{V_{11}^{2}V_{23}V_{33}}\\ \arg\frac{V_{12}V_{21}^{2}V_{32}}{V_{11}V_{22}^{2}V_{31}}&\arg\frac{V_{13}V_{22}^{2}V_{33}}{V_{12}V_{23}^{2}V_{32}}&\arg\frac{V_{11}V_{23}^{2}V_{31}}{V_{13}V_{21}^{2}V_{33}}\\ \arg\frac{V_{12}V_{22}V_{31}^{2}}{V_{11}V_{21}V_{32}^{2}}&\arg\frac{V_{13}V_{23}V_{32}^{2}}{V_{12}V_{22}V_{33}^{2}}&\arg\frac{V_{11}V_{21}V_{33}^{2}}{V_{13}V_{23}V_{31}^{2}}\\ \end{pmatrix}$ $=T^{2}\Phi T^{2}-T\Phi T^{2}\,.$ (37)
Alternatively, by multiplying $T$ from the right and using $T^{2}=T^{T}=T^{-1}$, it leads to a more compact notation
$\Delta T^{2}-\Delta T=T^{2}\Phi-T\Phi\,.$ (38)
This represents the transfer along the columns of the matrix elements,
$\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}\,.$ (39)
Here, the indices are taken modulo three. Note that when the transfer $T^{T}$ acts from the left, the indices of elements increase. These sum rules (39) have six independent constraints, because the relation for one of the columns is trivial. These relations are numerically verified as
$\begin{pmatrix}65.16^{\circ}&92.44^{\circ}&115.91^{\circ}\\ 114.89^{\circ}&157.51^{\circ}&1.089^{\circ}\\ 93.45^{\circ}&23.54^{\circ}&156.49^{\circ}\\ \end{pmatrix}-\begin{pmatrix}115.91^{\circ}&65.16^{\circ}&92.44^{\circ}\\ 1.089^{\circ}&114.89^{\circ}&157.51^{\circ}\\ 156.49^{\circ}&93.45^{\circ}&23.54^{\circ}\\ \end{pmatrix}=\begin{pmatrix}-50.76&27.29&23.47\\ 113.80&42.62&-156.42\\ -63.04&-69.91&132.95\\ \end{pmatrix}$ (40)
$=\begin{pmatrix}64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ 1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ \end{pmatrix}-\begin{pmatrix}114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ 1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ \end{pmatrix}$. (41)
Similarly, we can find a transfer along rows as
$T\Delta-\Delta=\begin{pmatrix}\arg\frac{V_{11}^{2}V_{32}V_{33}}{V_{12}V_{13}V_{31}^{2}}&\arg\frac{V_{12}^{2}V_{31}V_{33}}{V_{11}V_{13}V_{32}^{2}}&\arg\frac{V_{13}^{2}V_{31}V_{32}}{V_{11}V_{12}V_{33}^{2}}\\ \arg\frac{V_{12}V_{13}V_{21}^{2}}{V_{11}^{2}V_{22}V_{23}}&\arg\frac{V_{11}V_{13}V_{22}^{2}}{V_{12}^{2}V_{21}V_{23}}&\arg\frac{V_{11}V_{12}V_{23}^{2}}{V_{13}^{2}V_{21}V_{22}}\\ \arg\frac{V_{22}V_{23}V_{31}^{2}}{V_{21}^{2}V_{32}V_{33}}&\arg\frac{V_{21}V_{23}V_{32}^{2}}{V_{22}^{2}V_{31}V_{33}}&\arg\frac{V_{21}V_{22}V_{33}^{2}}{V_{23}^{2}V_{31}V_{32}}\\ \end{pmatrix}=T^{2}\Phi T^{2}-T^{2}\Phi T\,.$ (42)
Alternatively, by multiplying $T$ from the left, one obtains
$T^{2}\Delta-T\Delta=\Phi T^{2}-\Phi T\,$, (43)
and for matrix elements
$\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}\,.$ (44)
Iterating these two transfers, we can find the sum rules for all combinations between $\delta^{(\alpha i)}$ and $\delta^{(\beta j)}$.
Finally, we reproduce the results of the previous work. Since it suffices to examine the relation between $\delta^{(13)}$ and $\delta^{(11)}$, the case of $\alpha=1,i=2$ in Eq. (39) becomes
$\delta^{(11)}-\delta^{(13)}=\Phi_{2,2}-\Phi_{3,2}\,,~~~\pi-\delta_{\rm KM}-\delta_{\rm PDG}=\alpha-\gamma\,.$ (45)
Thus, the relation $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ is confirmed. Furthermore, it has also been shown in the previous work Yang:2025law ; Yang:2025ftl , such a difference between phases and angles is expressed in terms of a third-order invariant Luo:2025wio
$\delta_{\rm PDG}-\gamma$ $=\arg\left[-{V_{us}V_{cd}V_{tb}\over\det V_{\rm CKM}}\right]=\pi-\delta_{\rm KM}-\alpha\,.$ (46)
A comprehensive study of this kind of relation will be performed in the subsequent work.
The results generalize the previous sum rule to the nine CP phases in the different parameterizations and angles of unitarity triangles. Since the constraints from these sum rules provide twelve conditions for the nine phases and nine angles, six independent fundamental observables remain.
IV Summary
In this letter, we present rephasing invariant formulae $\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}\det V}\right]$ for CP phases $\delta^{(\alpha i)}$ associated with the nine parameterizations $V^{(\alpha i)}$ of a flavor mixing matrix. Here, $\alpha$ and $i$ denote the row and column with trivial phases in $V^{(\alpha i)}$.
Furthermore, we find exact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}$ and $\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ between the nine phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of the unitarity triangles. These twelve relations revealed intrinsic underlying orders connecting the CP phases and angles, which thought to be random and independent. We show that a relation from the previous work $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ is immediately reproduced as a special case ($\alpha=1,i=2$) of the first sum rules. These results provide a general and concise framework between differences of angles and phases, and offer a valuable guide for future experimental tests of CP violation.
Acknowledgment
The study is partly supported by the MEXT Leading Initiative for Excellent Young Researchers Grant Number JP2023L0013.

### v2

@@
-Rephasing invariant formulae for general parameterizations of flavor mixing matrix and exact sum rules with unitarity triangles
+Rephasing invariant formulae for CP phases in general parameterizations of flavor mixing matrix and exact sum rules with unitarity triangles
 Abstract
@@
 Here, $\alpha$ and $i$ denote the row and column carrying the trivial phases in a given parameterization.
-Furthermore, we show that the phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of unitarity triangles satisfy compact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}$ and $\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ where all indices are taken cyclically modulo three.
-These twelve relations are natural generalizations of the previous result $\delta_{\mathrm{PDG}}+\delta_{\mathrm{KM}}=\pi-\alpha+\gamma.$
+Furthermore, we show that the phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of unitarity triangles satisfy compact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha+1,i}-\Phi_{\alpha+2,i}$ and $\delta^{(\alpha+1,i)}-\delta^{(\alpha+2,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ where all indices are taken cyclically modulo three.
+These relations are natural generalizations of the previous result $\delta_{\mathrm{PDG}}+\delta_{\mathrm{KM}}=\pi-\alpha+\gamma.$
 preprint: STUPP-25-285
@@
 The flavor mixing of quarks and leptons has been one of the essential pieces for the development of particle physics, and it is also an intrinsically interesting subject of mathematical investigation.
-Among various representations of the mixing matrix Kobayashi:1973fv ; Chau:1984fp ; Hall:1993ni ; Kaur:2023ypg ; Yang:2020qsa ; Yang:2020goc ; Yang:2024ulq ; Yang:2025yst , nine Euler-angle-like parameterizations in terms of three rotation matrices and a distinct CP phase admit nine possible forms Fritzsch:1995nx ; Rasin:1997pn .
+Among various representations of the mixing matrix Kobayashi:1973fv ; Chau:1984fp ; Hall:1993ni ; Fritzsch:1995nx ; Kaur:2023ypg ; Yang:2020qsa ; Yang:2020goc ; Yang:2024ulq ; Yang:2025yst , Euler-angle-like parameterizations in terms of three rotation matrices and a distinct CP phase admit nine possible forms Rasin:1997pn .
 However, at first glance, these nine phases seem to take chaotic values, showing no apparent pattern.
-On the other hand, unitarity conditions of the mixing matrix give rise to six unitarity triangles Wu:1994di ; Xing:2009eg ; Harrison:2009bz ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Harrison:2025rkp and relations between the CP phases and the unitarity triangles has also been extensively studied Koide:2004gj ; Koide:2008yu ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2011bq ; Zhou:2011xm ; Qin:2011ub ; Qin:2010hn ; Zhang:2012ys ; Li:2012zxa ; Zhang:2012bk .
-These nine angles also seem to have independent values, showing no obvious structure.
+On the other hand, unitarity conditions of the mixing matrix give rise to six unitarity triangles Wu:1994di ; Xing:2009eg ; Harrison:2009bz ; He:2013rba ; He:2016dco ; Xing:2019tsn ; Harrison:2025rkp and a lot of papers have also been studied relations between the CP phases and the triangles Koide:2004gj ; Koide:2008yu ; Frampton:2010ii ; Dueck:2010fa ; Frampton:2010uq ; Li:2010ae ; Qin:2011bq ; Zhou:2011xm ; Qin:2011ub ; Qin:2010hn ; Zhang:2012ys ; Li:2012zxa ; Zhang:2012bk .
+These nine angles also seem to have independent values.
 More recently, rephasing invariant formulae for the CP phases Yang:2025hex ; Yang:2025law ; Yang:2025ftl produce a simple sum rule relating the CP phases and angles $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$.
-This led to a discovery of an underlying order between the random CP phases and angles.
-In this letter, to establish exact sum rules connecting the nine CP phases and the nine angles of unitarity triangles, we first derive rephasing invariant formulae for each CP phase in terms of elements of the mixing matrix $V$ and $\det V$.
-Next, we present sum rules between the phases and angles, showing that the previously known sum rule arises as a special case of these relations.
+This led to the discovery of an underlying order between the random CP phases and angles.
+In this letter, to organize relations connecting the nine CP phases and the nine angles of unitarity triangles, we first derive rephasing invariant formulae for each CP phase in terms of elements of the mixing matrix $V$ and $\det V$.
+Next, we present exact sum rules between the phases and angles, demonstrating that the previous sum rule is a special case of these relations.
 II Rephasing invariant formulae for CP phases in general parameterizations of flavor mixing matrix
@@
 $R_{12}(\theta)=\begin{pmatrix}c_{\theta}&s_{\theta}&0\\ -s_{\theta}&c_{\theta}&0\\ 0&0&1\\ \end{pmatrix},~~~R_{23}(\sigma)=\begin{pmatrix}1&0&0\cr 0&c_{\sigma}&s_{\sigma}\\ 0&-s_{\sigma}&c_{\sigma}\\ \end{pmatrix},~~R_{31}(\tau)=\begin{pmatrix}c_{\tau}&0&s_{\tau}\\ 0&1&0\\ -s_{\tau}&0&c_{\tau}\\ \end{pmatrix}$, (1)
-where $s_{\theta}=\sin\theta,c_{\theta}=\cos\theta.$ Furthermore, we define complex rotation matrices $R_{12}(\theta,\delta)$, $R_{23}(\sigma,\delta)$, and $R_{31}(\tau,\delta)$ by replacing $1\rightarrow e^{-i\delta}$ in Eq.
+where $s_{\theta}=\sin\theta,c_{\theta}=\cos\theta.$ Furthermore, complex rotation matrices $R_{12}(\theta,\delta)$, $R_{23}(\sigma,\delta)$, and $R_{31}(\tau,\delta)$ are defined by replacing $1\rightarrow e^{-i\delta}$ in Eq.
 (1).
-With these matrices, the mixing matrix $V$ admits nine different parameterizations Fritzsch:1997st
-$P1:~V^{(33)}$ $=R_{12}(\theta)R_{23}(\sigma,\delta^{(33)})R^{-1}_{12}(\theta^{\prime})\,,~~P2:~V^{(11)}=R_{23}(\sigma)R_{12}(\theta,\delta^{(11)})R^{-1}_{23}(\sigma^{\prime})\,$, (2)
-$P3:~V^{(13)}$ $=R_{23}(\sigma)R_{31}(\tau,\delta^{(13)})R_{12}(\theta)\,,~~~P4:~V^{(31)}=R_{12}(\theta)R_{31}(\tau,\delta^{(31)})R^{-1}_{23}(\sigma)\,$, (3)
-$P5:~V^{(22)}$ $=R_{31}(\tau)R_{12}(\theta,\delta^{(22)})R^{-1}_{31}(\tau^{\prime})\,,~~P6:~V^{(32)}=R_{12}(\theta)R_{23}(\sigma,\delta^{(32)})R_{31}(\tau)\,$, (4)
-$P7:~V^{(12)}$ $=R_{23}(\sigma)R_{12}(\theta,\delta^{(12)})R^{-1}_{31}(\tau)\,,~~P8:~V^{(21)}=R_{31}(\tau)R_{12}(\theta,\delta^{(21)})R_{23}(\sigma)\,$, (5)
+With these matrices, the flavor mixing matrix $V$ admits nine different parameterizations Fritzsch:1997st
+$P1:~V^{(33)}$ $=R_{12}(\theta)R_{23}(\sigma,\delta^{(33)})R^{-1}_{12}(\theta^{\prime})\,,~~P2:~V^{(11)}=R_{23}(\sigma)R_{12}(\theta,\delta^{(11)})R^{-1}_{23}(\sigma^{\prime})\,$,
+$P3:~V^{(13)}$ $=R_{23}(\sigma)R_{31}(\tau,\delta^{(13)})R_{12}(\theta)\,,~~~P4:~V^{(31)}=R_{12}(\theta)R_{31}(\tau,\delta^{(31)})R^{-1}_{23}(\sigma)\,$,
+$P5:~V^{(22)}$ $=R_{31}(\tau)R_{12}(\theta,\delta^{(22)})R^{-1}_{31}(\tau^{\prime})\,,~~P6:~V^{(32)}=R_{12}(\theta)R_{23}(\sigma,\delta^{(32)})R_{31}(\tau)\,$,
+$P7:~V^{(12)}$ $=R_{23}(\sigma)R_{12}(\theta,\delta^{(12)})R^{-1}_{31}(\tau)\,,~~P8:~V^{(21)}=R_{31}(\tau)R_{12}(\theta,\delta^{(21)})R_{23}(\sigma)\,$,
 $P9:~V^{(23)}$ $=R_{31}(\tau)R_{23}(\sigma,\delta^{(23)})R^{-1}_{12}(\theta)\,.$ (2)
@@
 Since there is a duplication for the element $V_{\alpha i}^{(\alpha i)}$, there are effectively five independent conditions, and the phase $\pi$ appears either zero or two times.
-Finally, if the Majorana phases are omitted,
+As the final condition, the argument of the determinant is
 $\arg\det V^{(\alpha i)}=-\delta^{(\alpha i)}\,.$ (4)
-In other words, the phase structure of these parameterizations $V^{(\alpha i)}$ is specified by these six conditions.
-Since the resulting CP phases are manifestly rephasing invariant, this construction remains valid even in the presence of Majorana phases for leptons.
-Let us consider transforming a mixing matrix $V$ in an arbitrary basis into one of these parameterizations $V^{(\alpha i)}$.
-Suppose that unphysical phases are removed by general phase transformations:
+In other words, the phase structures of these parameterizations $V^{(\alpha i)}$ are specified by these six conditions.
+Let us consider transforming the mixing matrix $V$ in an arbitrary basis into one of these parameterizations $V^{(\alpha i)}$.
+Suppose that general rephasing transformations remove unphysical phases as
 $V^{(\alpha i)}=\Psi_{L}^{\dagger}V\Psi_{R},$ (5)
@@
 $V=\Psi_{L}V^{(\alpha i)}\Psi_{R}^{\dagger}\,,~~V_{\beta j}=e^{i\gamma_{L\beta}}V^{(\alpha i)}_{\beta j}e^{-i\gamma_{Rj}}\,.$ (6)
-From a relation $\arg\det V=-\delta^{(\alpha i)}+\sum_{\beta,j}(\gamma_{L\beta}-\gamma_{Rj})$ obtained by taking the argument of determinant, the CP phase $\delta^{(\alpha i)}$ is found to be
-$\delta^{(\alpha i)}$ $=(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})-\arg\det V$ (11)
+By taking arguments of the determinants, a relation $\arg\det V=-\delta^{(\alpha i)}+\sum_{\beta,j}(\gamma_{L\beta}-\gamma_{Rj})$ is obtained and the CP phase $\delta^{(\alpha i)}$ is found to be
+$\delta^{(\alpha i)}$ $=(\gamma_{L1}+\gamma_{L2}+\gamma_{L3}-\gamma_{R1}-\gamma_{R2}-\gamma_{R3})-\arg\det V$
 $=\arg[V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}/V_{\alpha i}^{3}]-\arg\det V=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right].$ (7)
-In practice, since $V_{\alpha i}$ appears twice in the numerator, cancellations occur between the numerator and denominator:
-$\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right]=\arg\left[\prod_{\beta\neq\alpha,j\neq i}{V_{\alpha j}V_{\beta i}\over V_{\alpha i}\det V}\right].$ (13)
-Since six phases are required to eliminate unphysical phases $\gamma_{L\alpha}$ and $\gamma_{Ri}$, any other combinations have nontrivial phases of $V^{(\alpha i)}_{\beta j}$ explicitly.
+In practice, since $V_{\alpha i}$ appears twice in the numerator, cancellations occur between the numerator and denominator as
+$\delta^{(\alpha i)}=\arg\left[{V_{\alpha 1}V_{\alpha 2}V_{\alpha 3}V_{1i}V_{2i}V_{3i}\over V_{\alpha i}^{3}\det V}\right]=\arg\left[{(\prod_{j\neq i}V_{\alpha j})\,(\prod_{\beta\neq\alpha}V_{\beta i})\over V_{\alpha i}\det V}\right].$ (8)
+Since six phases are required to eliminate unphysical phases $\gamma_{L\alpha}$ and $\gamma_{Ri}$, any other combinations have nontrivial phases of $V^{(\alpha i)}_{\beta j}$ and are not solved for $\delta^{(\alpha i)}$ explicitly.
+This derivation remains valid even in the presence of Majorana phases for leptons because the resulting CP phases are manifestly rephasing invariant.
 Explicitly, the formulae are expressed as
-$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$ (14)
-$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$ (15)
+$\delta^{(11)}$ $=\arg\left[{V_{12}V_{13}V_{21}V_{31}\over V_{11}\det V}\right],~~\delta^{(12)}=\arg\left[{V_{11}V_{13}V_{22}V_{32}\over V_{12}\det V}\right],~~\delta^{(13)}=\arg\left[{V_{11}V_{12}V_{23}V_{33}\over V_{13}\det V}\right],$
+$\delta^{(21)}$ $=\arg\left[{V_{22}V_{23}V_{11}V_{31}\over V_{21}\det V}\right],~~\delta^{(22)}=\arg\left[{V_{21}V_{23}V_{12}V_{32}\over V_{22}\det V}\right],~~\delta^{(23)}=\arg\left[{V_{21}V_{22}V_{13}V_{33}\over V_{23}\det V}\right],$
 $\delta^{(31)}$ $=\arg\left[{V_{32}V_{33}V_{11}V_{21}\over V_{31}\det V}\right],~~\delta^{(32)}=\arg\left[{V_{31}V_{33}V_{12}V_{22}\over V_{32}\det V}\right],~~\delta^{(33)}=\arg\left[{V_{31}V_{32}V_{13}V_{23}\over V_{33}\det V}\right].$ (9)
-They provide a necessary and sufficient set of “fifth-order” invariants, because only nine such invariants can be constructed using $\det V$.
-By using the CKM matrix parameters from the latest UTfit results UTfit:2022hsi
-$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$, (17)
+They provide a necessary and sufficient set of “irreducible fifth-order” invariants which cannot be decomposed into second- and third-orders, because only nine such invariants are constructed using $\det V$.
+From the latest UTfit results of the CKM matrix parameters UTfit:2022hsi
+$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~~~\sin\theta_{23}=0.04200\pm 0.00047\,$,
 $\sin\theta_{13}$ $=0.003714\pm 0.000092\,,~~~\delta=1.137\pm 0.022=65.15^{\circ}\pm 1.3^{\circ}\,$, (10)
 the numerical evaluation at the best-fit values yields
-$\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix}=\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.16^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}$.
+$\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix}=\begin{pmatrix}92.44^{\circ}&115.91^{\circ}&65.15^{\circ}\\ 157.51^{\circ}&1.089^{\circ}&114.89^{\circ}\\ 23.54^{\circ}&156.49^{\circ}&93.45^{\circ}\\ \end{pmatrix}$.
 (11)
@@
 $\delta_{\rm PDG}=\delta^{(13)}\,,~~~\delta_{\rm KM}=\pi-\delta^{(11)}\,.$ (12)
-Interestingly, a sum over any row or column of the phases $\delta^{(ij)}$ yields the same value
+Interestingly, a sum over any row or column of the phases $\delta^{(ij)}$ yields the same value of a ninth-order invariant
 $\sum_{i}\delta^{(ij)}=\sum_{j}\delta^{(ij)}=\arg\left[{\prod_{\alpha,i}V_{\alpha i}\over\det V^{3}}\right]=273.49^{\circ}\,.$ (13)
@@
 These values agree with values obtained from moduli of matrix elements,
-$\cos\delta^{(11)}$ $=\frac{\left(1-|V_{11}|^{2}\right)^{2}|V_{22}|^{2}-|V_{11}V_{12}V_{21}|^{2}-|V_{13}V_{31}|^{2}}{2|V_{11}V_{12}V_{13}V_{21}V_{31}|}=-0.0425\,$, (23)
-$\cos\delta^{(12)}$ $=\frac{\left(1-|V_{12}|^{2}\right)^{2}|V_{23}|^{2}-|V_{12}V_{13}V_{22}|^{2}-|V_{11}V_{32}|^{2}}{2|V_{11}V_{12}V_{13}V_{22}V_{32}|}=-0.437\,$, (24)
-$\cos\delta^{(13)}$ $=-\frac{\left(1-|V_{13}|^{2}\right)^{2}|V_{22}|^{2}-|V_{12}V_{23}V_{13}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{12}V_{13}V_{23}V_{33}|}=0.420\,$, (25)
-$\cos\delta^{(21)}$ $=-\frac{\left(1-|V_{21}|^{2}\right)^{2}|V_{12}|^{2}-|V_{11}V_{21}V_{22}|^{2}-|V_{23}V_{31}|^{2}}{2|V_{21}V_{22}V_{23}V_{11}V_{31}|}=-0.924\,$, (26)
-$\cos\delta^{(22)}$ $=\frac{\left(1-|V_{22}|^{2}\right)^{2}|V_{33}|^{2}-|V_{22}V_{23}V_{32}|^{2}-|V_{12}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{12}V_{32}|}=0.9998\,$, (27)
-$\cos\delta^{(23)}$ $=-\frac{\left(1-|V_{23}|^{2}\right)^{2}|V_{32}|^{2}-|V_{22}V_{23}V_{33}|^{2}-|V_{13}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{13}V_{33}|}=-0.421\,$, (28)
-$\cos\delta^{(31)}$ $=-\frac{\left(1-|V_{31}|^{2}\right)^{2}|V_{22}|^{2}-|V_{21}V_{32}V_{31}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{21}V_{31}V_{32}V_{33}|}=0.917\,$, (29)
-$\cos\delta^{(32)}$ $=-\frac{\left(1-|V_{32}|^{2}\right)^{2}|V_{23}|^{2}-|V_{22}V_{32}V_{33}|^{2}-|V_{12}V_{31}|^{2}}{2|V_{31}V_{32}V_{33}V_{12}V_{22}|}=-0.917\,$, (30)
+$\cos\delta^{(11)}$ $=\frac{\left(1-|V_{11}|^{2}\right)^{2}|V_{22}|^{2}-|V_{11}V_{12}V_{21}|^{2}-|V_{13}V_{31}|^{2}}{2|V_{11}V_{12}V_{13}V_{21}V_{31}|}=-0.0425\,$,
+$\cos\delta^{(12)}$ $=\frac{\left(1-|V_{12}|^{2}\right)^{2}|V_{23}|^{2}-|V_{12}V_{13}V_{22}|^{2}-|V_{11}V_{32}|^{2}}{2|V_{11}V_{12}V_{13}V_{22}V_{32}|}=-0.437\,$,
+$\cos\delta^{(13)}$ $=-\frac{\left(1-|V_{13}|^{2}\right)^{2}|V_{22}|^{2}-|V_{12}V_{23}V_{13}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{12}V_{13}V_{23}V_{33}|}=0.420\,$,
+$\cos\delta^{(21)}$ $=-\frac{\left(1-|V_{21}|^{2}\right)^{2}|V_{12}|^{2}-|V_{11}V_{21}V_{22}|^{2}-|V_{23}V_{31}|^{2}}{2|V_{21}V_{22}V_{23}V_{11}V_{31}|}=-0.924\,$,
+$\cos\delta^{(22)}$ $=\frac{\left(1-|V_{22}|^{2}\right)^{2}|V_{33}|^{2}-|V_{22}V_{23}V_{32}|^{2}-|V_{12}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{12}V_{32}|}=0.9998\,$,
+$\cos\delta^{(23)}$ $=-\frac{\left(1-|V_{23}|^{2}\right)^{2}|V_{32}|^{2}-|V_{22}V_{23}V_{33}|^{2}-|V_{13}V_{21}|^{2}}{2|V_{21}V_{22}V_{23}V_{13}V_{33}|}=-0.421\,$,
+$\cos\delta^{(31)}$ $=-\frac{\left(1-|V_{31}|^{2}\right)^{2}|V_{22}|^{2}-|V_{21}V_{32}V_{31}|^{2}-|V_{11}V_{33}|^{2}}{2|V_{11}V_{21}V_{31}V_{32}V_{33}|}=0.917\,$,
+$\cos\delta^{(32)}$ $=-\frac{\left(1-|V_{32}|^{2}\right)^{2}|V_{23}|^{2}-|V_{22}V_{32}V_{33}|^{2}-|V_{12}V_{31}|^{2}}{2|V_{31}V_{32}V_{33}V_{12}V_{22}|}=-0.917\,$,
 $\cos\delta^{(33)}$ $=\frac{\left(1-|V_{33}|^{2}\right)^{2}|V_{11}|^{2}-|V_{13}V_{31}V_{33}|^{2}-|V_{23}V_{32}|^{2}}{2|V_{13}V_{31}V_{33}V_{23}V_{32}|}=-0.0603\,.$ (15)
-Since $\delta^{(\alpha i)}$ also carries the sign of $\sin\delta^{(\alpha i)}$, it contains more information than an evaluation based on absolute values of matrix elements.
-III Sum rules between CP phases in general parameterizations and unitarity triangles
-It is noteworthy to revisit relations between the results and angles of unitarity triangles. The three angles of the unitarity triangle are rephasing invariants,
+Since $\delta^{(\alpha i)}$ also carries the sign of $\sin\delta^{(\alpha i)}$, it contains more information than such an evaluation based on absolute values of matrix elements.
+III Exact sum rules between CP phases in general parameterizations and unitarity triangles
+It is noteworthy to reconsider relations between the results and angles of unitarity triangles. The three angles of the unitarity triangle are rephasing invariants,
 $\alpha=\arg\left[-{V_{td}V_{tb}^{*}\over V_{ud}V_{ub}^{*}}\right]=92.40^{\circ},~~\beta=\arg\left[-{V_{cd}V_{cb}^{*}\over V_{td}V_{tb}^{*}}\right]=22.49^{\circ},~~\gamma=\arg\left[-{V_{ud}V_{ub}^{*}\over V_{cd}V_{cb}^{*}}\right]=65.11^{\circ}.$ (16)
@@
 $\sum_{i}\Phi_{ij}=\sum_{j}\Phi_{ij}=\pi\,.$ (19)
-Since solving general sum rules is a difficult task, we employ a transfer-matrix approach. Defining a phase matrix $\Delta$ and a transfer matrix $T$,
+Since solving general sum rules is a challenging task, we employ a transfer matrix approach. Defining a phase matrix $\Delta$ and a transfer matrix $T$,
 $\Delta\equiv\begin{pmatrix}\delta^{(11)}&\delta^{(12)}&\delta^{(13)}\\ \delta^{(21)}&\delta^{(22)}&\delta^{(23)}\\ \delta^{(31)}&\delta^{(32)}&\delta^{(33)}\end{pmatrix},~~T\equiv\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\\ \end{pmatrix}$, (20)
@@
 $\Delta T^{2}-\Delta T=T^{2}\Phi-T\Phi\,.$ (22)
-This represents the transfer along the columns of the matrix elements,
-$\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}\,.$ (39)
-Here, the indices are taken modulo three.
-Note that when the transfer $T^{T}$ acts from the left, the indices of elements increase.
-These sum rules (39) have six independent constraints, because the relation for one of the columns is trivial.
-These relations are numerically verified as
-$\begin{pmatrix}65.16^{\circ}&92.44^{\circ}&115.91^{\circ}\\ 114.89^{\circ}&157.51^{\circ}&1.089^{\circ}\\ 93.45^{\circ}&23.54^{\circ}&156.49^{\circ}\\ \end{pmatrix}-\begin{pmatrix}115.91^{\circ}&65.16^{\circ}&92.44^{\circ}\\ 1.089^{\circ}&114.89^{\circ}&157.51^{\circ}\\ 156.49^{\circ}&93.45^{\circ}&23.54^{\circ}\\ \end{pmatrix}=\begin{pmatrix}-50.76&27.29&23.47\\ 113.80&42.62&-156.42\\ -63.04&-69.91&132.95\\ \end{pmatrix}$ (40)
+This represents the transfer along columns of $\Delta$ and rows of $\Phi$;
+$\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha+1,i}-\Phi_{\alpha+2,i}\,.$ (23)
+Here, the indices are taken modulo three. Note that the transfer $T^{T}$ from the left increases the indices of rows. These relations are numerically verified as
+$\begin{pmatrix}65.15^{\circ}&92.44^{\circ}&115.91^{\circ}\\ 114.89^{\circ}&157.51^{\circ}&1.089^{\circ}\\ 93.45^{\circ}&23.54^{\circ}&156.49^{\circ}\\ \end{pmatrix}-\begin{pmatrix}115.91^{\circ}&65.15^{\circ}&92.44^{\circ}\\ 1.089^{\circ}&114.89^{\circ}&157.51^{\circ}\\ 156.49^{\circ}&93.45^{\circ}&23.54^{\circ}\\ \end{pmatrix}=\begin{pmatrix}-50.76&27.29&23.47\\ 113.80&42.62&-156.42\\ -63.04&-69.91&132.95\\ \end{pmatrix}$
 $=\begin{pmatrix}64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ 114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ 1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ \end{pmatrix}-\begin{pmatrix}114.85^{\circ}&65.11^{\circ}&0.0370^{\circ}\\ 1.054^{\circ}&22.49^{\circ}&156.46^{\circ}\\ 64.09^{\circ}&92.40^{\circ}&23.50^{\circ}\\ \end{pmatrix}$.
 (24)
-Similarly, we can find a transfer along rows as
+Similarly, we can find another transfer as
 $T\Delta-\Delta=\begin{pmatrix}\arg\frac{V_{11}^{2}V_{32}V_{33}}{V_{12}V_{13}V_{31}^{2}}&\arg\frac{V_{12}^{2}V_{31}V_{33}}{V_{11}V_{13}V_{32}^{2}}&\arg\frac{V_{13}^{2}V_{31}V_{32}}{V_{11}V_{12}V_{33}^{2}}\\ \arg\frac{V_{12}V_{13}V_{21}^{2}}{V_{11}^{2}V_{22}V_{23}}&\arg\frac{V_{11}V_{13}V_{22}^{2}}{V_{12}^{2}V_{21}V_{23}}&\arg\frac{V_{11}V_{12}V_{23}^{2}}{V_{13}^{2}V_{21}V_{22}}\\ \arg\frac{V_{22}V_{23}V_{31}^{2}}{V_{21}^{2}V_{32}V_{33}}&\arg\frac{V_{21}V_{23}V_{32}^{2}}{V_{22}^{2}V_{31}V_{33}}&\arg\frac{V_{21}V_{22}V_{33}^{2}}{V_{23}^{2}V_{31}V_{32}}\\ \end{pmatrix}=T^{2}\Phi T^{2}-T^{2}\Phi T\,.$ (25)
@@
 and for matrix elements
-$\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}\,.$ (44)
+$\delta^{(\alpha+1,i)}-\delta^{(\alpha+2,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}\,.$ (27)
 Iterating these two transfers, we can find the sum rules for all combinations between $\delta^{(\alpha i)}$ and $\delta^{(\beta j)}$.
 Finally, we reproduce the results of the previous work. Since it suffices to examine the relation between $\delta^{(13)}$ and $\delta^{(11)}$, the case of $\alpha=1,i=2$ in Eq. (23) becomes
-$\delta^{(11)}-\delta^{(13)}=\Phi_{2,2}-\Phi_{3,2}\,,~~~\pi-\delta_{\rm KM}-\delta_{\rm PDG}=\alpha-\gamma\,.$ (45)
+$\delta^{(11)}-\delta^{(13)}=\Phi_{22}-\Phi_{32}\,,~~~\pi-\delta_{\rm KM}-\delta_{\rm PDG}=\alpha-\gamma\,.$ (28)
 Thus, the relation $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ is confirmed.
@@
 The results generalize the previous sum rule to the nine CP phases in the different parameterizations and angles of unitarity triangles.
-Since the constraints from these sum rules provide twelve conditions for the nine phases and nine angles, six independent fundamental observables remain.
 IV Summary
@@
 Here, $\alpha$ and $i$ denote the row and column with trivial phases in $V^{(\alpha i)}$.
-Furthermore, we find exact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha-2,i}-\Phi_{\alpha-1,i}$ and $\delta^{(\alpha-2,i)}-\delta^{(\alpha-1,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ between the nine phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of the unitarity triangles.
-These twelve relations revealed intrinsic underlying orders connecting the CP phases and angles, which thought to be random and independent.
+In practice, since cancellations occur between the numerator and the denominator, these are a necessary and sufficient set of irreducible fifth-order invariants constructed using the determinant.
+Furthermore, we find exact sum rules $\delta^{(\alpha,i+2)}-\delta^{(\alpha,i+1)}=\Phi_{\alpha+1,i}-\Phi_{\alpha+2,i}$ and $\delta^{(\alpha+1,i)}-\delta^{(\alpha+2,i)}=\Phi_{\alpha,i+2}-\Phi_{\alpha,i+1}$ between the nine phases $\delta^{(\alpha i)}$ and the nine angles $\Phi_{\alpha i}$ of the unitarity triangles.
+These relations revealed intrinsic underlying orders connecting the CP phases and angles, which seem to be random and independent.
 We show that a relation from the previous work $\delta_{\rm PDG}+\delta_{\rm KM}=\pi-\alpha+\gamma$ is immediately reproduced as a special case ($\alpha=1,i=2$) of the first sum rules.
-These results provide a general and concise framework between differences of angles and phases, and offer a valuable guide for future experimental tests of CP violation.
+These results provide a general and concise framework for differences between angles and phases, and offer a valuable guide for future experimental tests of CP violation.
 Acknowledgment